OpenStudy (anonymous):

how do i find domain of y=ln(3-2x)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

myininaya (myininaya):

3-2x>0 will give you your domain 3-2x>0 means you want the inside to be positive ln(negative) does not exist ln(zero) does not exist

myininaya (myininaya):

3>2x 3/2>x x<3/2 (-infinity,3/2)

OpenStudy (anonymous):

so ln doesnt matter..

myininaya (myininaya):

yes it does matter we want the inside of natural to be positive is it something else natural log will not exist

myininaya (myininaya):

if we just have the polynomail, y=3-2x then the domain is all real numbers x can be anything there is a output for every input if we have y=ln(x), then x>0 if we have y=ln(3-2x), then 3-2x>0

OpenStudy (anonymous):

ohhh thats how we know its not negative kay thank yu

myininaya (myininaya):

yes